OpenStudy (anonymous):

What are the zeros of the polynomial function: f(x) = x^3 – x^2 - 12x ?

OpenStudy (anonymous):

\(x=0\) is one by inspection

OpenStudy (anonymous):

so -4 and 3 would be the others?

OpenStudy (anonymous):

so you can factor as \(f(x)=x(x^2-x-12)\) and find the other two by factoring

OpenStudy (anonymous):

= x(x^2 - x - 12) x= 0

OpenStudy (anonymous):

i think you might have the signs wrong, lets check

OpenStudy (anonymous):

yea and 2 more - as satellite indicated

OpenStudy (anonymous):